# A Unified Framework for Variable Selection in Model-Based Clustering with Missing Not at Random

### v2

@@ 2 Background
Model-based Clustering and Parsimonious Mixture Form. Model-based clustering adopts a probabilistic framework by assuming that the data matrix ${\mathbf{Y}}=({\bm{y}}_{1},\dots,{\bm{y}}_{N})^{\top}\in\mathds{R}^{N\times D}$, with each ${\bm{y}}_{n}\in\mathds{R}^{D}$, is independently drawn from a location-scale mixture distribution, a class known for its universal approximation capabilities and favorable convergence properties [25, 55, 14, 53, 52, 46, 47, 49, 60, 26, 27]. The density of an observation ${\bm{y}}_{n}$ under a $K$-component GMM with parameters ${\bm{\alpha}}=\{{\bm{\pi}},{\bm{\mu}},{\bm{\Sigma}}\}$ is given by: $f_{\text{GMM}}({\bm{y}}_{n}\mid K,m,{\bm{\alpha}})=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}({\bm{y}}_{n}\mid{\bm{\mu}}_{k},{\bm{\Sigma}}_{k,m})$ where $\pi_{k}>0$ and $\sum_{k=1}^{K}\pi_{k}=1$. The term $\mathcal{N}({\bm{y}}_{n}\mid{\bm{\mu}}_{k},{\bm{\Sigma}}_{k,m})$ denotes the multivariate normal density with mean vector ${\bm{\mu}}_{k}$ and covariance matrix ${\bm{\Sigma}}_{k,m}$, which is compactly denoted by ${\bm{\Theta}}_{k,m}=({\bm{\mu}}_{k},{\bm{\Sigma}}_{k,m})$. The covariance matrix ${\bm{\Sigma}}_{k,m}$ encodes the structure determined by model form $m$, allowing for parsimonious modeling via spectral decompositions that control cluster volume, shape, and orientation [12]. These constraints are particularly valuable in high-dimensional settings, where $D\gg N$ and standard GMMs require estimating ${\mathcal{O}}(K^{2}D)$ parameters, leading to overparameterization. Model selection criteria such as the BIC [58, 39, 22, 21, 48], slope heuristics [6, 51, 50], integrated classification likelihood (ICL) [7, 24], extended BIC (eBIC) [23, 45], dendrogram selection criterion [17, 69], and Sin-White information criterion (SWIC) [61, 72] can be employed to select both the number of components $K$ and the covariance structure $m$.
@@
Penalized Log-Likelihood Methods for Simultaneous Clustering and Variable Selection. We follow the framework of [11], which extends [75, 13], introducing cluster-specific penalties via group-wise weighting matrices $\bf\pi_{k}$ for adaptive regularization in Gaussian graphical mixture models. This method improves upon stepwise procedures [38, 37] by handling high-dimensional data more efficiently. The penalized log-likelihood is given by:
@@ 3 Our Proposal: Variable Selection in Model-Based Clustering with MNAR
$f({\bm{Y}}^{o}\mid K,m,r,l,{\mathbb{V}},{\bm{\Theta}})=\prod_{n=1}^{N}\left(\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}\left({\bm{y}}_{n}^{o}\mid\tilde{{\bm{\nu}}}_{k,o},\tilde{\Delta}_{k,oo}\right)\right),$ (4)
where $\tilde{{\bm{\nu}}}_{k,o}=\begin{pmatrix}{\bm{\nu}}_{k,o}\\ {\bm{\gamma}}_{o}\end{pmatrix},\quad\tilde{{\bm{\Delta}}}_{k,oo}=\begin{pmatrix}{\bm{\Delta}}_{k,oo}&\mathbf{0}\\ \mathbf{0}&{\bm{\Gamma}}_{oo}\end{pmatrix}$. Here, ${\bm{\nu}}_{k,o},{\bm{\gamma}}_{o}$ and ${\bm{\Delta}}_{k,oo},{\bm{\Delta}}_{k,oo}$ correspond to the block means and covariances from the $\mathbb{S}$$\mathbb{R}$} model. The derivation extends the technical proof from [39] and is detailed in the supplement. This representation offers: (i) a unified GMM encoding the variable roles in ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$ via observed-data parameters, (ii) compatibility with standard EM algorithms, enabling MAR-aware estimation and internal imputation that respects cluster structure, and (iii) theoretical guarantees of identifiability and consistency in selecting the true variable partition via Equation˜2 (see Section 4).
Incorporating MNARz Mechanism into ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$. To explicitly address MNAR data, we integrate the MNARz mechanism from [62], where missingness depends on the latent class membership. Estimation proceeds via EM, while identifiability and the transformation of MNARz into MAR on augmented data (i.e., original data with the missingness indicator matrix $\mathbf{C}$) are retained (see Section 4). Let ${\bm{y}}_{n}=({\bm{y}}_{n}^{{\mathbb{S}}},{\bm{y}}_{n}^{{\mathbb{U}}},{\bm{y}}_{n}^{{\mathbb{W}}})$ denote the full data for observation $n$, partitioned by role (Section 2), and let ${\bm{c}}_{n}$ be its binary missingness pattern. With $z_{n}\in[K]$ denoting the latent class, the complete-data density under $z_{nk}=1$ is
@@
$\ell_{\text{comp}}({\bm{\Theta}};{\bm{Y}},\mathbf{C})=\sum_{n=1}^{N}\sum_{k=1}^{K}z_{nk}\log\bigl(\pi_{k}f_{k}({\bm{y}}_{n};{\bm{\alpha}}_{k},\bm{\xi})f_{c}({\bm{c}}_{n};{\bm{\psi}}_{k})\bigr).$ (6)
@@
$\ell({\bm{\Theta}};{\bm{Y}},\mathbf{C})=\sum_{n=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}\,f_{c}^{\text{MNARz}}({\bm{c}}_{n,M^{\prime}};{\bm{\psi}}_{k})f_{k,M}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\psi}}_{M})\right].$ (7)
@@
$\ell({\bm{\Theta}};{\bm{Y}}^{o},\mathbf{C})=\sum_{n=1}^{N}\log\sum_{k=1}^{K}\pi_{k}f_{c}({\bm{c}}_{n};{\bm{\psi}}_{k})f_{k}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\gamma}}),$
where $f_{k}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\gamma}}):=\int f_{\text{clust}}({\bm{y}}_{n}^{S}\mid{\bm{\alpha}}_{k})f_{\text{reg}}({\bm{y}}_{n}^{U}\mid{\bm{y}}_{n}^{R};{\bm{\theta}}_{\text{reg}})f_{\text{indep}}({\bm{y}}_{n}^{W}\mid{\bm{\theta}}_{\text{indep}})\,d{\bm{y}}_{n}^{m}$.
@@
Practical Initialization. Like all mixture EM algorithms, ours is sensitive to initialization. We use a robust warm start: (i) fast single imputation $\tilde{{\bm{Y}}}$ (only for ranking); (ii) $K$-means++/hierarchical clustering/random start on $\tilde{{\bm{Y}}}$ for $(\pi_{k},\mu_{k},\Sigma_{k})$; (iii) given initial partitions from step (ii) $\bm{\Psi}_{k}^{(0)}$ set to diagonal estimates using samples assigned to each cluster (cluster-aware estimates); (iv) $\psi_{k}$ initialized from class-wise missing rates after a soft E-step. We then run a $\lambda$-path (as illustrated below) for Stage A, followed by unpenalized SRUW estimation for Stage B. Per-$K$ penalty grids are constructed by data-dependent upper bounds and a geometric path: For $\lambda$, with the hard partition ${\bm{Z}}^{(0)}$ from initialization,
@@
followed by the same $L$-point geometric path $\rho_{\ell}=\rho_{\max}(\rho_{\min}/\rho_{\max})^{(\ell-1)/(L-1)}$ with $\rho_{\min}=\xi\rho_{\max}$. We set $\operatorname{diag}({\bm{P}}_{k})=\mathbf{0}$ when we do not impose penalty on diagonal entries.
@@ 5 Experiments
Simulated Dataset. We generate two synthetic data sets with sample size $n=2000$ for variable/model selection with missing values treatment benchmarking. The first data set mimics [39] where each observation ${\bm{y}}_{n}\in\mathds{R}^{7}$ is drawn from a four-component Gaussian mixture on the clustering block ${\mathbb{S}}=\{{\bm{y}}_{1},{\bm{y}}_{2},{\bm{y}}_{3}\}$. The components are equiprobable and have means $(0,0,0)^{{}^{\prime}},\,(-6,6,0)^{{}^{\prime}},\,(0,0,6)^{{}^{\prime}},\,(-6,6,6)^{{}^{\prime}}$ with a diagonal covariance matrix ${\bm{\Sigma}}=\text{diag}(6\sqrt{2},1,2)\,\sigma_{\text{scale}}^{2}$. The redundant block ${\mathbb{U}}=\{{\bm{y}}_{4},{\bm{y}}_{5}\}$ is generated by a linear regression ${\mathbb{U}}=(-1,2)^{{}^{\prime}}+\mathbf{X}_{S}[(0.5,2)^{{}^{\prime}},(1,0)^{{}^{\prime}}]+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}\bigl(\bf 0,\mathbf{R}(\pi/6)\text{diag}(1,3)\mathbf{R}(\pi/6)^{{}^{\prime}}\bigr)$ and $\mathbf{R}({\bm{\theta}})$ denotes the usual planar rotation. The remaining $W$ variables are iid $\mathcal{N}(0,1)$ noise. In the second design an observation ${\bm{y}}_{n}\in\mathds{R}^{14}$. The clustering block now consists of $({\bm{y}}_{1},{\bm{y}}_{2})$, drawn from an equiprobable four-component Gaussian mixture with means $(0,0),(4,0),(0,2),(4,2)$ and common covariance $0.5\,\bf I_{2}$. Conditional on $({\bm{y}}_{1},{\bm{y}}_{2})$, the vector ${\bm{y}}_{3{:}14}$ follows a linear model ${\bm{y}}_{3{:}14}={\bm{\alpha}}+({\bm{y}}_{1},{\bm{y}}_{2})\bm{\beta}+\bm{\varepsilon}$, where $\bm{\alpha}$}, the $2\times 12$ coefficient matrix $\bm{\beta}$ and the diagonal covariance $\bm{\Omega}$} are varied across eight sub-scenarios: scenario $1$ contains only pure noise, scenarios $2\text{-}7$ introduce more regressors, and scenario $8$ adds intercept shifts together with three additional pure-noise variables ${\bm{y}}_{12{:}14}$. In our setting, we consider $K\in\{2,3,4\}$ and every method except VarSelLCM is allowed to learn its own covariance structure. We vary the missing ratio $\{0.05,0.1,0.2,0.3,0.5\}$ under both MAR and MNAR patterns. For second simulated dataset, we temporarily consider only scenario $8$. Moreover, for Selvar and SelvarMNARz, we fix $c=2$ and compute ${\bm{P}}_{k}$ with the spectral distance. Further details of the synthetic-data generation scheme are provided in the supplement. From Figure 1, we observe that SelvarMNARz delivers the highest ARI and the lowest WNRMSE across every missing-data level in both scenarios. Its performance declines only modestly as missingness increases, while the impute-then-cluster baselines deteriorate sharply, especially at 30% and 50% missingness under MNAR, pointing to the benefit of modeling the missing-data mechanism within the clustering algorithm. To statistically validate that our model exhibits a significantly slower performance decline under high missingness, we conducted one-sided Welch’s t-tests (Bonferroni-corrected) at the 50% MNAR level. As shown in Table 1, the ARI of SelvarMNARz is significantly greater than all baseline models ($p<0.001$) across 20 replications.
@@
Transcriptome Data. We re-examine the 1,267-gene Arabidopsis thaliana transcriptome dataset previously analyzed using SelvarClust and SelvarClustMV in [37, 39] to illustrate clustering with variable selection. Full experimental details are provided in the Supplementary Material; here, we summarize the settings and key findings. We fitted SelvarMNARz for $K\in\{2,\dots,20\}$ clusters with $c=5$, spectral distance weights ${\bm{P}}_{k}$, and a $\pi_{k}LC$ mixture structure as in [39]. For conciseness, we refer to Project 1 as P1, Project 2 as P2, and so on. As shown in Figure 3, the proposed framework partitions the transcriptome into 18 clusters and reaffirms P1-P4 as core drivers, while globally assigning P5-P7 as being explained by these drivers. This finding is consistent with [39], which also identifies P2 as relevant for clustering, supporting the hypothesis that iron signaling responses are critical for defining co-expression groups when a broader gene set (including entries with missing data) is considered. Additionally, all three methods agree that P5 is not a primary clustering factor and is more likely a correlated effect explained by broader cellular processes. However, a key departure lies in the reclassification of P6 and P7, previously identified as informative clustering drivers [37, 39], into a redundant $\mathbb{U}$} block, explained by P1-P4. This suggests that, after accounting for MNAR effects, variation in P6-P7 is largely predictable from the core axes P1-P4, and P5-P7 all measure late-stage stress outputs. Further analysis of cluster sizes and the per-cluster $R^{2}$ values, shown in Table 2, validates the selected variables when regressing P5-P7 on P1-P4 within each cluster. Six biologically coherent clusters (6, 7, 8, 10, 12, 18) exhibit $R^{2}>0.60$, indicating that late-stress variation is well encoded by P1-P4. Cluster 1, although large, is transcriptionally flat across both $\mathbb{S}$} and $\mathbb{U}$}; its low $R^{2}=0.02$ thus reflects a vanishing denominator rather than hidden structure. True residual structure (unexplained $\mathbb{U}$} signal) is concentrated in two groups: clusters 5 and 11, which are exactly where the 18th mixture component forms

### v3

@@ 2 Background
Model-based Clustering and Parsimonious Mixture Form. Model-based clustering adopts a probabilistic framework by assuming that the data matrix ${\mathbf{Y}}=({\bm{y}}_{1},\dots,{\bm{y}}_{N})^{\top}\in\mathds{R}^{N\times D}$, with each ${\bm{y}}_{n}\in\mathds{R}^{D}$, is independently drawn from a location-scale mixture distribution, a class known for its universal approximation capabilities and favorable convergence properties [25, 55, 14, 53, 52, 46, 47, 49, 60, 26, 27]. The density of an observation ${\bm{y}}_{n}$ under a $K$-component GMM with parameters ${\bm{\alpha}}=\{{\bm{\pi}},{\bm{\mu}},{\bm{\Sigma}}\}$ is given by: $f_{\text{GMM}}({\bm{y}}_{n}\mid K,m,{\bm{\alpha}})=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}({\bm{y}}_{n}\mid{\bm{\mu}}_{k},{\bm{\Sigma}}_{k})$ where $\pi_{k}>0$ and $\sum_{k=1}^{K}\pi_{k}=1$. The term $\mathcal{N}({\bm{y}}_{n}\mid{\bm{\mu}}_{k},{\bm{\Sigma}}_{k})$ denotes the multivariate normal density with mean vector ${\bm{\mu}}_{k}$ and covariance matrix ${\bm{\Sigma}}_{k}$, which is compactly denoted by ${\bm{\Theta}}_{k}=({\bm{\mu}}_{k},{\bm{\Sigma}}_{k})$. The covariance matrix ${\bm{\Sigma}}_{k}$ encodes the structure determined by model form $m$, allowing for parsimonious modeling via spectral decompositions that control cluster volume, shape, and orientation [12]. These constraints are particularly valuable in high-dimensional settings, where $D\gg N$ and standard GMMs require estimating ${\mathcal{O}}(K^{2}D)$ parameters, leading to overparameterization. Model selection criteria such as the BIC [58, 39, 22, 21, 48], slope heuristics [6, 51, 50], integrated classification likelihood (ICL) [7, 24], extended BIC (eBIC) [23, 45], dendrogram selection criterion [17, 69], and Sin-White information criterion (SWIC) [61, 72] can be employed to select both the number of components $K$ and the covariance structure $m$.
@@
Penalized Log-Likelihood Methods for Simultaneous Clustering and Variable Selection. We follow the framework of [11], which extends [75, 13], introducing cluster-specific penalties via group-wise weighting matrices ${\bm{P}}_{k}$ for adaptive regularization in Gaussian graphical mixture models. This method improves upon stepwise procedures [38, 37] by handling high-dimensional data more efficiently. The penalized log-likelihood is given by:
@@ 3 Our Proposal: Variable Selection in Model-Based Clustering with MNAR
$f({\bm{Y}}^{o}\mid K,m,r,l,{\mathbb{V}},{\bm{\Theta}})=\prod_{n=1}^{N}\left(\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}\left({\bm{y}}_{n}^{o}\mid\tilde{{\bm{\nu}}}_{k,o},\tilde{{\bm{\Delta}}}_{k,oo}\right)\right),$ (4)
where $\tilde{{\bm{\nu}}}_{k,o}=\begin{pmatrix}{\bm{\nu}}_{k,o}\\ {\bm{\gamma}}_{o}\end{pmatrix},\quad\tilde{{\bm{\Delta}}}_{k,oo}=\begin{pmatrix}{\bm{\Delta}}_{k,oo}&\mathbf{0}\\ \mathbf{0}&{\bm{\Gamma}}_{oo}\end{pmatrix}$. Here, ${\bm{\nu}}_{k,o},{\bm{\gamma}}_{o}$ and ${\bm{\Delta}}_{k,oo},{\bm{\Gamma}}_{oo}$ correspond to the block means and covariances from the $\mathbb{S}$$\mathbb{R}$} model. The derivation extends the technical proof from [39] and is detailed in the supplement. This representation offers: (i) a unified GMM encoding the variable roles in ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$ via observed-data parameters, (ii) compatibility with standard EM algorithms, enabling MAR-aware estimation and internal imputation that respects cluster structure, and (iii) theoretical guarantees of identifiability and consistency in selecting the true variable partition via Equation˜2 (see Section 4).
Incorporating MNARz Mechanism into ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$. To explicitly address MNAR data, we integrate the MNARz mechanism from [62], where missingness depends on the latent class membership. Estimation proceeds via EM, while identifiability and the transformation of MNARz into MAR on augmented data (i.e., original data with the missingness indicator matrix $\bm{C}$}) are retained (see Section 4). Let ${\bm{y}}_{n}=({\bm{y}}_{n}^{{\mathbb{S}}},{\bm{y}}_{n}^{{\mathbb{U}}},{\bm{y}}_{n}^{{\mathbb{W}}})$ denote the full data for observation $n$, partitioned by role (Section 2), and let ${\bm{c}}_{n}$ be its binary missingness pattern. With $z_{n}\in[K]$ denoting the latent class, the complete-data density under $z_{nk}=1$ is
@@
$\ell_{\text{comp}}({\bm{\Theta}};{\bm{Y}},{\bm{C}})=\sum_{n=1}^{N}\sum_{k=1}^{K}z_{nk}\log\bigl(\pi_{k}f_{k}({\bm{y}}_{n};{\bm{\alpha}}_{k},{\bm{\xi}})f_{c}({\bm{c}}_{n};{\bm{\psi}}_{k})\bigr).$ (6)
@@
$\ell({\bm{\Theta}};{\bm{Y}},{\bm{C}})=\sum_{n=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}\,f_{c}^{\text{MNARz}}({\bm{c}}_{n,M^{\prime}};{\bm{\psi}}_{k})f_{k,M}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\psi}}_{M})\right].$ (7)
@@
$\ell({\bm{\Theta}};{\bm{Y}}^{o},{\bm{C}})=\sum_{n=1}^{N}\log\sum_{k=1}^{K}\pi_{k}f_{c}({\bm{c}}_{n};{\bm{\psi}}_{k})f_{k}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\xi}}),$
where $f_{k}^{o}({\bm{y}}_{n}^{o};{\bm{\alpha}}_{k},{\bm{\xi}}):=\int f_{\text{clust}}({\bm{y}}_{n}^{S}\mid{\bm{\alpha}}_{k})f_{\text{reg}}({\bm{y}}_{n}^{U}\mid{\bm{y}}_{n}^{R};{\bm{\theta}}_{\text{reg}})f_{\text{indep}}({\bm{y}}_{n}^{W}\mid{\bm{\theta}}_{\text{indep}})\,d{\bm{y}}_{n}^{m}$.
@@
Practical Initialization. Like all mixture EM algorithms, ours is sensitive to initialization. We use a robust warm start: (i) fast single imputation $\tilde{{\bm{Y}}}$ (only for ranking); (ii) $K$-means++/hierarchical clustering/random start on $\tilde{{\bm{Y}}}$ for $(\pi_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}_{k})$; (iii) given initial partitions from step (ii) $\bm{\Psi}_{k}^{(0)}$ set to diagonal estimates using samples assigned to each cluster (cluster-aware estimates); (iv) ${\bm{\psi}}_{k}$ initialized from class-wise missing rates after a soft E-step. We then run a $\lambda$-path (as illustrated below) for Stage A, followed by unpenalized SRUW estimation for Stage B. Per-$K$ penalty grids are constructed by data-dependent upper bounds and a geometric path: For $\lambda$, with the hard partition ${\bm{Z}}^{(0)}$ from initialization,
@@
followed by the same $L$-point geometric path $\rho_{\ell}=\rho_{\max}(\rho_{\min}/\rho_{\max})^{(\ell-1)/(L-1)}$ with $\rho_{\min}=\xi\rho_{\max}$. We set $\operatorname{diag}({\bm{P}}_{k})=\mathbf{0}$ when we do not impose a penalty on diagonal entries.
@@ 5 Experiments
Simulated Dataset. We generate two synthetic data sets with sample size $n=2000$ for variable/model selection with missing values treatment benchmarking. The first data set mimics [39] where each observation ${\bm{y}}_{n}\in\mathds{R}^{7}$ is drawn from a four-component Gaussian mixture on the clustering block ${\mathbb{S}}=\{{\bm{y}}_{1},{\bm{y}}_{2},{\bm{y}}_{3}\}$. The components are equiprobable and have means $(0,0,0)^{{}^{\prime}},\,(-6,6,0)^{{}^{\prime}},\,(0,0,6)^{{}^{\prime}},\,(-6,6,6)^{{}^{\prime}}$ with a diagonal covariance matrix ${\bm{\Sigma}}=\text{diag}(6\sqrt{2},1,2)\,\sigma_{\text{scale}}^{2}$. The redundant block ${\mathbb{U}}=\{{\bm{y}}_{4},{\bm{y}}_{5}\}$ is generated by a linear regression ${\mathbb{U}}=(-1,2)^{{}^{\prime}}+\mathbf{X}_{S}[(0.5,2)^{{}^{\prime}},(1,0)^{{}^{\prime}}]+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}\bigl(\bf 0,\mathbf{R}(\pi/6)\text{diag}(1,3)\mathbf{R}(\pi/6)^{{}^{\prime}}\bigr)$ and $\mathbf{R}({\bm{\theta}})$ denotes the usual planar rotation. The remaining $W$ variables are iid $\mathcal{N}(0,1)$ noise. In the second design, an observation ${\bm{y}}_{n}\in\mathds{R}^{14}$. The clustering block now consists of $({\bm{y}}_{1},{\bm{y}}_{2})$, drawn from an equiprobable four-component Gaussian mixture with means $(0,0),(4,0),(0,2),(4,2)$ and common covariance $0.5\,\bf I_{2}$. Conditional on $({\bm{y}}_{1},{\bm{y}}_{2})$, the vector ${\bm{y}}_{3{:}14}$ follows a linear model ${\bm{y}}_{3{:}14}={\bm{\alpha}}+({\bm{y}}_{1},{\bm{y}}_{2})\bm{\beta}+\bm{\varepsilon}$, where $\bm{\alpha}$}, the $2\times 12$ coefficient matrix $\bm{\beta}$ and the diagonal covariance $\bm{\Omega}$} are varied across eight sub-scenarios: scenario $1$ contains only pure noise, scenarios $2\text{-}7$ introduce more regressors, and scenario $8$ adds intercept shifts together with three additional pure-noise variables ${\bm{y}}_{12{:}14}$. In our setting, we consider $K\in\{2,3,4\}$ and every method except VarSelLCM is allowed to learn its own covariance structure. We vary the missing ratio $\{0.05,0.1,0.2,0.3,0.5\}$ under both MAR and MNAR patterns. For the second simulated dataset, we temporarily consider only scenario $8$. Moreover, for Selvar and SelvarMNARz, we fix $c=2$ and compute ${\bm{P}}_{k}$ with the spectral distance. Further details of the synthetic-data generation scheme are provided in the supplement. From Figure 1, we observe that SelvarMNARz delivers the highest ARI and the lowest WNRMSE across every missing-data level in both scenarios. Its performance declines only modestly as missingness increases, while the impute-then-cluster baselines deteriorate sharply, especially at 30% and 50% missingness under MNAR, pointing to the benefit of modeling the missing-data mechanism within the clustering algorithm. To statistically validate that our model exhibits a significantly slower performance decline under high missingness, we conducted one-sided Welch’s t-tests (Bonferroni-corrected) at the 50% MNAR level. As shown in Table 1, the ARI of SelvarMNARz is significantly greater than all baseline models ($p<0.001$) across 20 replications.
@@
Transcriptome Data. We re-examine the 1,267-gene Arabidopsis thaliana transcriptome dataset previously analyzed using SelvarClust and SelvarClustMV in [37, 39] to illustrate clustering with variable selection. Full experimental details are provided in the Supplementary Material; here, we summarize the settings and key findings. We fitted SelvarMNARz for $K\in\{2,\dots,20\}$ clusters with $c=5$, spectral distance weights ${\bm{P}}_{k}$, and a $\pi_{k}LC$ mixture structure as in [39]. For conciseness, we refer to Project 1 as P1, Project 2 as P2, and so on. As shown in Figure 3, the proposed framework partitions the transcriptome into 18 clusters and reaffirms P1-P4 as core drivers, while globally assigning P5-P7 as being explained by these drivers. This finding is consistent with [39], which also identifies P2 as relevant for clustering, supporting the hypothesis that iron signaling responses are critical for defining co-expression groups when a broader gene set (including entries with missing data) is considered. Additionally, all three methods agree that P5 is not a primary clustering factor and is more likely a correlated effect explained by broader cellular processes. However, a key departure lies in the reclassification of P6 and P7, previously identified as informative clustering drivers [37, 39], into a redundant $\mathbb{U}$} block, explained by P1-P4. This suggests that, after accounting for MNAR effects, variation in P6-P7 is largely predictable from the core axes P1-P4, and P5-P7 all measure late-stage stress outputs. Further analysis of cluster sizes and the per-cluster $R^{2}$ values, shown in Table 2, validates the selected variables when regressing P5-P7 on P1-P4 within each cluster. Six biologically coherent clusters (6, 7, 8, 10, 12, 18) exhibit $R^{2}>0.60$, indicating that late-stress variation is well encoded by P1-P4. Cluster 1, although large, is transcriptionally flat across both $\mathbb{S}$} and $\mathbb{U}$}; its low $R^{2}=0.02$ thus reflects a vanishing denominator rather than hidden structure. True residual structure (unexplained $\mathbb{U}$} signal) is concentrated in two groups: clusters 5 and 11, which are exactly where the 18th mixture component forms.